OpenStudy (anonymous):

what is the right endpoint for y=2sin(x-pi/4)

OpenStudy (anonymous):

@myininaya can you help me with this?

OpenStudy (anonymous):

@robtobey i need help with this

OpenStudy (anonymous):

@kc_kennylau can you help with this

OpenStudy (anonymous):

@wolfe8 can you help with this?

OpenStudy (anonymous):

i have the amp,per, and le now i just need the re @kc_kennylau

OpenStudy (kc_kennylau):

what's the le

OpenStudy (anonymous):

left endpoint @kc_kennylau

OpenStudy (kc_kennylau):

i mean the value @shanmv98

OpenStudy (anonymous):

oh pi/4 @kc_kennylau

OpenStudy (kc_kennylau):

And what's its period?

OpenStudy (kc_kennylau):

Add the left endpoint and the period together, and you'll get the right endpoint.

OpenStudy (anonymous):

so is it 3pi/4 @kc_kennylau

OpenStudy (kc_kennylau):

nope, what's its period?

OpenStudy (anonymous):

the period is 2pi @kc_kennylau

OpenStudy (kc_kennylau):

how did you calculate 2pi?

OpenStudy (anonymous):

i just knew that the period is usually 2pi @kc_kennylau

OpenStudy (kc_kennylau):

oh yep it's 2pi, then add pi/4 with 2pi what do you get?

OpenStudy (anonymous):

i dont know how to do it @kc_kennylau

OpenStudy (kc_kennylau):

\[\frac\pi4+2\pi=?\]

OpenStudy (kc_kennylau):

C'mon this is just basic stuff

OpenStudy (anonymous):

9pi/4 @kc_kennylau

OpenStudy (kc_kennylau):

exactly :D